OpenStudy (anonymous):

How to study for the AP Calc AB test?!?!?! HELP OMG this is so panic-making...

OpenStudy (anonymous):

Open book, read.

OpenStudy (primeralph):

Open stuff; read stuff.

OpenStudy (anonymous):

TT0TT any fast way to review?

ganeshie8 (ganeshie8):

I think your notes is the best thing to stick to when reviewing... touching new resources can make u loose time however u can practice new problems if you have already finished with your textbook and others.... just my general opinion...

OpenStudy (anonymous):

Also, youtube is great :) in 10 minutes sometimes you can learn a whole unit!

OpenStudy (mathmale):

That assumes you've already read the assigned sections of the book and need only a review. I sense you need more than that. All I can suggest at this point is to pick out some practice problems, post them here one at a time, and ask questions to help yourself get started out solving those problems. Can you do that? Please at least try one problem now.

OpenStudy (anonymous):

I'm doing a practice AP Calc AB test in class and it's KILLING ME SO UTTERLY...and my teacher says that it's for a grade, and I can't solve any of them all of the way through!!!!!!!!

OpenStudy (mathmale):

Lucy, I truly sympathize, but where have you been? The time to start preparing for an exam is well before the test. Have you done problems similar to this practice AP test before? I hope so. If not, why not?

OpenStudy (anonymous):

I haven't done problems like these...and it's because I had no idea they would be this hard...! I know the fundamental stuff in general and I have a high B in the class, yet I have no idea what to do~ TT0TT

OpenStudy (mathmale):

So you've learned something truly valuable: Start early and judge how much time you're going to need to complete the more challenging questions. Again, if you want help now, choose and post one question at a time, and do at least something, and share it, before asking for help. Congrats on your high B and your knowing the "fundamental stuff." Let's now put that knowledge to good use.

OpenStudy (anonymous):

TToTT

OpenStudy (anonymous):

uhm ok so how do you integrate a(t)=[4t-3]/[e^4t-3] ? It's part of a bigger word problem, and I feel that this is the first step but I don't know quite where to start...

OpenStudy (anonymous):

hi ganeshie8 <3

ganeshie8 (ganeshie8):

I saw you doing hard differential equations and integration problems many weeks ago. Actually you know a lot and very good for AP test... just need more practice i feel...

OpenStudy (beccaboo333):

Aw Ganesh baby <3 someone else doing hearts for you :3 Study, have people test you. Create questions and answer them to train your brain. Good luck.

OpenStudy (anonymous):

more practice...ok~ will do! haha first i have to complete this AP test my teacher let us take home... it's due tomorrow and I have 5 question s left...the others I kind of made an educated guess on heetee~

ganeshie8 (ganeshie8):

\(\large \int \frac{4t-3}{e^{4t-3}}dt\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

and I'm having trouble with generally integrating e raised to any function...such as tanx...

ganeshie8 (ganeshie8):

okay, looks "integration by parts" is your weakness

Miracrown (miracrown):

Well there is no really fast way, you have to put in the time to prepare to the exam, there are no shortcuts to it unfortunately It is like this for many high level courses. Do you have a particular problem on what you're currently struggling with? We can help you with that @lucy4104

OpenStudy (anonymous):

If you have a solution book, you can always do review sections of that and if you're not getting it the solution books usually show the steps. Also, sometimes teachers put up practice tests and put "last years test" and put very difficult questions on it at times even though it wasn't the actual test from the year prior, this is to scare students. And your integral looks like a u sub + by parts.

OpenStudy (mathmale):

In this case you'd need u=a substation meant to simplify the appearance of the integral, along with du=the differential of u. Supposing you wanted to integrate:\[\int\limits_{}^{}e ^{\tan x}dx\] ... as stated, that'd be tough. However, supposing you let u=tan x and also have \[du=\sec ^{2}x dx\] then the integration would be relatively easy:\[\int\limits_{}^{}e ^{\tan x}\sec ^{2}x*dx\] This would be simplifiable to \[\int\limits_{}^{}e^u*du=e^u + C.\] What would the last step be?

OpenStudy (mathmale):

Again, Lucy, do as many actual problems from that practice exam as you can. Start with the easier ones so that you can build up confidence in y our ability to solve such problems.

OpenStudy (anonymous):

@Miracrown uhm, integrating a(t)=[4t-3]/[e^4t-3], and in general integrating e raised to a function~ as well as this one other problem: A spherical balloon is being blown up. When its volum is 36(pi) cm^3, and increasing at the rate 20(pi) cm^2/sec, how fast is its diamete changing at that moment in cm/sec?

OpenStudy (anonymous):

@mathmale wait hold on, Are you saying that I can put in another variable??????

OpenStudy (mathmale):

That's a "related rates" problem. Share what you've done towards solving this problem. Someone will then give you feedback.

OpenStudy (anonymous):

well, it's still x, but still...!

ganeshie8 (ganeshie8):

one problem at a time, finish with tanx first..

OpenStudy (mathmale):

Yes, but you're introducing time by finding the derivative of x with respect to time. Thanks, @ganeshie8: sound advice.

OpenStudy (anonymous):

ok e^tanx, then the 4t-3 thing, then the related rates

ganeshie8 (ganeshie8):

you cant integrate e^tanx using elementary methods

ganeshie8 (ganeshie8):

what mathmale cooked up is an easy problem : \(\large \int e^{\tan x} \sec^2x dx\)

ganeshie8 (ganeshie8):

go thru it again quick

OpenStudy (anonymous):

hmmm I know that e^x*dx rule but I'm still not quite sure what it's telling me to do...

OpenStudy (anonymous):

would that come out as just e^tanx / [secx]?

Miracrown (miracrown):

Referring to your problem, is that e to the power of 4t, or to the power of 4t - 3? @lucy4104

Miracrown (miracrown):

in your power, I'm asking you to what power is the e taken to What is it's exponent? Is it just 4t, or 4t - 3? I can't tell through here is that -3 is part of the exponent, or just subtracting to the side?

ganeshie8 (ganeshie8):

\(\large a(t) = \frac{4t-3}{e^{4t-3}}\)

ganeshie8 (ganeshie8):

^^ @Miracrown

OpenStudy (anonymous):

yesh, that thing ^^

ganeshie8 (ganeshie8):

we need to use "integration by parts"

Miracrown (miracrown):

Okay, so it is an exponent. Alright, so any ideas how we might approach this problem?

OpenStudy (anonymous):

ue^tanx or the 4t-3 one?

OpenStudy (anonymous):

srry diregrd that u in da front

Miracrown (miracrown):

So, we are going to have to use an integration technique here, do you remember what the two main techniques are?

Miracrown (miracrown):

Just the general integration techniques, there are two special ones...

Miracrown (miracrown):

One starts with a u

OpenStudy (anonymous):

uhm...one is most likeyly going to be the (int sign) e^u du= e^u, and then the other is maybe...natural log?

Miracrown (miracrown):

Well that just shows what the integration of e^u is, what I was asking as in regards to integration techniques, not just specific to e. The two I am referring to are u substitution, and integration by parts. Do those sound familiar?

OpenStudy (anonymous):

u substitution sounds familiar...integration by parts, not so much -_-""

Miracrown (miracrown):

Alright, well for this we will be using u substitution Do you remember how it starts?

OpenStudy (anonymous):

would u substute 4t-3 for u?

Miracrown (miracrown):

Very good, yes!

Miracrown (miracrown):

What would we do next?

Miracrown (miracrown):

Need help?

OpenStudy (anonymous):

maybe...hehe...yeah

OpenStudy (anonymous):

so integrating u/e^u

Miracrown (miracrown):

Ok, Well remember this is a u substitution So we have to find du How can we find du?

OpenStudy (anonymous):

du is 4~

Miracrown (miracrown):

good

Miracrown (miracrown):

Well not just that We do replace the 4t - 3 with u But notice that du = 4*dt What does dt equal?

OpenStudy (anonymous):

du/4

Miracrown (miracrown):

Right, so we replace dt with du/4 - And the 1/4 can be moved outside

OpenStudy (anonymous):

so...put a 1/4 inside, and a 4 on the outside?

Miracrown (miracrown):

No just the 1/4 outside, we moved the 1/4 from the du, to the outside Since we have (1/4)*du

OpenStudy (anonymous):

oh, it just tags along?oh okay~ so the 1/4 in and 4 out is only if ur working with the raw problem, not using u?

Miracrown (miracrown):

The 1/4 is from replacing dt with du/4 The 1/4 is from replacing dt with du/4 outside

OpenStudy (anonymous):

yeah, so it tags along with the du and then since its a constant, you can move that out...so the 1/4 in and 4 out is only if ur working with the raw problem, not using u...right...? idkidk

Miracrown (miracrown):

I'm not sure where you are getting there is a 4 from There is just a 1/4, no 4

Miracrown (miracrown):

^^

Miracrown (miracrown):

Yes, you can move a constant out, which is what was done, we moved the 1/4 from du/4 outside

OpenStudy (anonymous):

ok so now we have [1/4](int sign) u/e^u du...what next....?

Miracrown (miracrown):

What might we be able to do next to integrate this?

Miracrown (miracrown):

Take a guess, Is there a way we can move the e^u to the top?

OpenStudy (anonymous):

yes, e^-u

Miracrown (miracrown):

There you go :) So we have this!

Miracrown (miracrown):